OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, 1). Which equation defines circle P?

OpenStudy (anonymous):

the equation of a circle in standard form is (x-h)^2 + (y-k)^2 = r^2 where (h,k) is your center and radius of r.

OpenStudy (anonymous):

just plug those values into the equation and you have it....

OpenStudy (anonymous):

Um what is H and K?

OpenStudy (anonymous):

h is the x coordinate of your center k is the y coordinate of your center

OpenStudy (anonymous):

they're given to you... and so is the radius r.

OpenStudy (anonymous):

oh sorry, So would the answer be (x+2)^2 + (y+1)^2=64?

OpenStudy (anonymous):

almost... but not quite... try again...

OpenStudy (anonymous):

btw, that one that you just posted is a circle centered at (-2, -1) and radius 8

OpenStudy (anonymous):

so it is (x – 2)^2 + (y – 1)^2 = 64 then? I think i see how i messed up.

OpenStudy (anonymous):

yep... good work man... :)

OpenStudy (anonymous):

Woo! Thank you show much for your help. I understand it now.

OpenStudy (anonymous):

yw